# LLM-Based Emulation of the Radio Resource Control Layer: Towards AI-Native RAN Protocols

### v4

@@ Abstract
Integrating Large AI Models (LAMs) into 6G mobile networks is a key enabler of the AI‑Native Air Interface (AI-AI), where protocol intelligence must scale beyond handcrafted logic. This paper presents, to our knowledge, the first standards-compliant emulation of the Radio Resource Control (RRC) layer using a decoder-only LAM (LLaMA-class) fine-tuned with Low‑Rank Adaptation (LoRA) on a multi-vendor corpus of real-world traces spanning both 5G and 4G systems. We treat RRC as a domain-specific language and construct a segmentation-safe, question–answer (Question-and-Answer (QA)) dataset that preserves Abstract Syntax Notation (ASN.1) structure through linearization prior to Byte Pair Encoding (BPE) tokenization. The proposed approach combines parameter-efficient adaptation with schema-bounded prompting to ensure syntactic and procedural fidelity. Evaluation introduces a standards-aware triad—ASN.1 conformance, field-level coverage analysis, and uplink-to-downlink state-machine checks—alongside semantic similarity and latency profiling across 120 configurations. On 30k 5G request–response pairs plus an additional 4.8k QA turns from 4G sessions, our 8B model achieves a median cosine similarity of 0.97, a 61% relative gain over a zero-shot baseline, while sustaining high conformance rates. These results demonstrate that LAMs, when augmented with protocol-aware reasoning, can directly orchestrate control-plane procedures, laying the foundation for the future Artificial Intelligence (AI)-native Radio Access Network (RAN).
@@ I Introduction
Large AI Models (LAMs) are poised to transform wireless networks by enabling autonomous, cognitive capabilities that go beyond the scope of traditional optimization methods[3]. The forthcoming sixth generation (6G) of mobile networks is expected to incorporate Artificial Intelligence (AI) methods not merely as auxiliary optimisation tools but also as integral design primitives of the AI‑Native Air Interface (AI-AI). Recent proposals for an AI-AI posit that Machine Learning (ML) models could co‑design Physical Layer (PHY) and Medium Access Control (MAC) procedures [15], thereby enabling radio stacks that adapt automatically to dynamic service requirements, resource constraints, and propagation conditions [4]. The rapid evolution of LAMs [10, 13] strengthens this vision—when endowed with multi-modal perception and reasoning, such models become plausible agents for protocol‑centric tasks that demand both linguistic competence and structured decision making. This raises a timely research challenge: determining the extent to which LAMs can directly engage in the generation and interpretation of control signaling in the Radio Access Network (RAN). Can LAMs achieve sufficient protocol literacy for RRC [1] emulation? What fine-tuning strategies are effective? What is the performance fidelity of such a system? These are the research questions this paper intends to answer. We hypothesize that LAM‑scale pretraining provides the compositional priors needed to generalize across 3GPP releases and vendor variants, and to perform multi‑procedure, stateful reasoning under incomplete context—capabilities that compact task‑specific transformers or grammar‑only decoders lack without extensive bespoke engineering.
One particularly compelling target for AI intervention is the Radio Resource Control (RRC) layer of 5G New Radio (NR). RRC is the Layer 3 (L3) of the cellular network. It orchestrates connection establishment, mobility and configuration between User Equipment (UE) and the Next‑Generation Node B (gNB). Conceptually, RRC messages form a specialized, domain‑specific language; learning this language is analogous to learning syntax and semantics in natural‑language processing. Treating RRC messages as such positions LAM ‑based sequence models as natural emulators of the RRC layer. Demonstrating accurate RRC emulation would constitute an early, concrete instantiation of AI-AI principles, with direct implications for the design of the 6G control plane.
Can the protocol literacy of a LAM —specifically, its ability to parse and generate 3GPP-compliant RRC messages—be systematically improved, and through which fine‑tuning strategies? This paper answers this question through a novel two‑phase methodology that (i) imparts RRC knowledge to a decoder‑only LAM via supervised fine‑tuning enhanced with Low‑Rank Adaptation (LoRA), and (ii) evaluates the resulting model’s fidelity in emulating RRC procedures under realistic network scenarios. The findings offer evidence about the feasibility of LAM ‑based standards-compliant RRC emulation.
@@ III Fine-tuning of LLM for RRC Message Processing
This section describes the pipeline for adapting a decoder-only LLM to emulate RRC procedures. Our approach consists of three components: (i) dataset construction, where raw traces are reorganized into segmentation-safe uplink–downlink Question-and-Answer (QA) pairs; (ii) parameter-efficient supervised fine-tuning using LoRA; and (iii) latency-aware evaluation to assess feasibility under control-plane timing constraints. Earlier traces often exceeded context limits, requiring ad-hoc segmentation that broke procedural continuity. Our design eliminates this by enforcing per-procedure boundaries.
@@ III-B Supervised Fine-Tuning Framework
where $\theta$ are trainable parameters, $y_{t}$ is the token generated at step $t$, and $T$ is the sequence length. This formulation aligns with the QA-style mapping introduced in Sec. III-A.
@@ III-D Decoding regimes
Unless otherwise stated, all reported results in Sec. IV use RRC_constrain for tuned configurations, as it delivers near-ceiling ASN.1/State-Machine Conformance (SMC) pass rates while maintaining practical latency.
@@ IV-C Latency and inference throughput
Weight-only INT4 (Q4_K_M) lowers medians by approximately 20–30% relative to FP16 without altering the decoding regime. For example, an 8B model under RRC_constrain drops from $\sim$3.6 s to $\sim$2.6 s, while a 3B backbone falls from $\sim$2.3 s to $\sim$1.7 s (Table V). Validator success remains high, with only minor reductions in ASN.1/SMC pass rates. Figure 8 visualizes the latency–fidelity frontier for all LTE configurations. The fastest points correspond to origin models, which achieve low latency by producing short, often invalid RRC messages (reflected in their poor semantic similarity). Conversely, tuned configurations deliver near-ceiling protocol validity (Tables VI–VII) but remain well above RRC timer budgets, with the best case at $\sim$1.6 s. Notably, smaller backbones (1B/3B) reduce latency substantially without compromising semantic fidelity, suggesting that backbone scaling combined with quantization and constrained decoding is a promising direction for future sub-100 ms targets.
@@ V Conclusion and Future Work
    Inference acceleration: speculative multi-token decoding [7], hierarchical agent architectures, and lightweight backbones to reconcile fidelity with real-time constraints.
@@
    Extended-context modeling: the context window length of current LLM s is critical for cross-layer reasoning and complex RRC procedures. Longer contexts or memory-augmented transformers (e.g., RoPE-extended variants, ALiBi [11]) will be explored to eliminate segmentation and preserve procedural continuity.

### v5

@@ Abstract
Integrating Large AI Models (LAMs) into 6G mobile networks is a key enabler of the AI‑Native Air Interface (AI-AI), where protocol intelligence must scale beyond handcrafted logic. This paper presents, to our knowledge, the first standards-compliant emulation of the Radio Resource Control (RRC) layer using a decoder-only LAM (LLaMA-class) fine-tuned with Low‑Rank Adaptation (LoRA) on a multi-vendor corpus of real-world traces spanning both 5G and 4G systems. We treat RRC as a domain-specific language and construct a segmentation-safe Question-and-Answer (QA) dataset that preserves Abstract Syntax Notation (ASN.1) structure through linearization prior to Byte Pair Encoding (BPE) tokenization. The proposed approach combines parameter-efficient adaptation with schema-bounded prompting to ensure syntactic and procedural fidelity. Evaluation introduces a standards-aware triad—ASN.1 conformance, field-level coverage analysis, and uplink-to-downlink state-machine checks—alongside semantic similarity and latency profiling across 120 configurations. On 30k 5G request–response pairs plus an additional 4.8k QA turns from 4G sessions, our 8B model achieves a median cosine similarity of 0.97, a 61% relative gain over a zero-shot baseline, while sustaining high conformance rates. These results demonstrate that LAMs, when augmented with protocol-aware reasoning, can directly orchestrate control-plane procedures, laying the foundation for the future Artificial Intelligence (AI)-native Radio Access Network (RAN).
@@ I Introduction
Large AI Models (LAMs) are poised to transform wireless networks by enabling autonomous, cognitive capabilities that go beyond the scope of traditional optimization methods[5]. The forthcoming sixth generation (6G) of mobile networks is expected to incorporate Artificial Intelligence (AI) methods not merely as auxiliary optimisation tools but also as integral design primitives of the AI‑Native Air Interface (AI-AI). Recent proposals for an AI-AI posit that Machine Learning (ML) models could co‑design Physical Layer (PHY) and Medium Access Control (MAC) procedures [20], thereby enabling radio stacks that adapt automatically to dynamic service requirements, resource constraints, and propagation conditions [7]. The rapid evolution of LAMs [13, 18] strengthens this vision—when endowed with multi-modal perception and reasoning, such models become plausible agents for protocol‑centric tasks that demand both linguistic competence and structured decision making. This raises a timely research challenge: determining the extent to which LAMs can directly engage in the generation and interpretation of control signaling in the Radio Access Network (RAN).
[…]
The 5G RRC layer orchestrates connection establishment, mobility and configuration between User Equipment (UE) and the Next‑Generation Node B (gNB). Conceptually, RRC messages form a specialized, domain‑specific language; learning this language is analogous to learning syntax and semantics in natural‑language processing. Treating RRC messages as such positions LAM ‑based sequence models as natural emulators of the RRC layer. Demonstrating accurate RRC emulation would constitute an early, concrete instantiation of AI-AI principles, with direct implications for the design of the 6G control plane.
Can the protocol literacy of a LAM —specifically, its ability to parse and generate 3GPP-compliant RRC messages—be systematically improved, and through which fine‑tuning strategies? This paper answers this question through a novel two‑phase methodology that (i) imparts RRC knowledge to a decoder‑only LAM via supervised fine‑tuning enhanced with LoRA, and (ii) evaluates the resulting model’s fidelity in emulating RRC procedures under realistic network scenarios. The findings offer evidence about the feasibility of LAM ‑based standards-compliant RRC emulation.
@@ III Fine-tuning of LLM for RRC Message Processing
This section describes the pipeline for adapting a decoder-only LLM to emulate RRC procedures. Our approach consists of three components: (i) dataset construction, where raw traces are reorganized into segmentation-safe uplink–downlink QA pairs; (ii) parameter-efficient supervised fine-tuning using LoRA; and (iii) latency-aware evaluation to assess feasibility under control-plane timing constraints. Earlier traces often exceeded context limits, requiring ad-hoc segmentation that broke procedural continuity. Our design eliminates this by enforcing per-procedure boundaries.
@@ III-B Supervised Fine-Tuning Framework
where $\theta$ are trainable parameters, $y_{t}$ is the token generated at step $t$, and $T$ is the sequence length. This formulation aligns with the QA-style mapping introduced in Sec. III-A. $X$ and $Y$ are tokenized using the same Byte Pair Encoding (BPE) tokenizer as the Llama backbone. Therefore, $T$ denotes the number of BPE tokens in the response sequence.
@@ III-C Training Configuration and Convergence Behavior
We note that earlier iterations of this training pipeline experienced frequent interruptions (e.g., SSD stalls, NCCL watchdog timeouts). These were identified as infrastructure-specific bottlenecks (insufficient storage I/O and driver-level software bugs) rather than model instability. Following the migration to the updated cluster (Table XI), the training process proved robust, achieving convergence without any hardware-induced interruptions.
@@ III-D Decoding regimes
Unless otherwise stated, all reported results in Sec. IV use RRC_constrain for tuned configurations, as it delivers near-ceiling ASN.1/State-Machine Conformance (SMC) pass rates while maintaining practical latency. Crucially, this schema-bounded approach mitigates the structural violations observed in unconstrained baselines (e.g., the duplication of fields). By explicitly injecting the session-specific ASN.1 definitions into the context, we leverage the model’s in-context learning capabilities to ground the generation process. This ensures that the model attends strictly to the fields declared in the active micro-schema, effectively preventing the hallucination of extraneous or repetitive elements without requiring computationally expensive token-level grammar constraints.
@@ IV-C Latency and inference throughput
Weight-only INT4 (Q4_K_M) lowers medians by approximately 20–30% relative to FP16 without altering the decoding regime. For example, an 8B model under RRC_constrain drops from $\sim$3.6 s to $\sim$2.6 s, while a 3B backbone falls from $\sim$2.3 s to $\sim$1.7 s (Table V). Validator success remains high, with only minor reductions in ASN.1/SMC pass rates. Figure 8 visualizes the latency–fidelity frontier for all LTE configurations. The fastest points correspond to origin models, which achieve low latency by producing short, often invalid RRC messages (reflected in their poor semantic similarity). Conversely, tuned configurations deliver near-ceiling protocol validity (Tables VI–VII) but remain well above RRC timer budgets, with the best case at $\sim$1.6 s. The latency gap between the “origin” and fine-tuned configurations (e.g., $\approx$ 338 ms vs. $\approx$ 2.2 s for Llama-3.2 1B in LABEL:tab:new_benchmark_agg_full) reveals an intrinsic trade-off between protocol completeness and generation time. The origin models achieve lower latency primarily because they fail to generate complex, lengthy message bodies (e.g., full RRCReconfiguration structures), resulting in significantly lower semantic similarity. In contrast, the fine-tuned models correctly produce these long sequences, linearly increasing inference time due to the autoregressive nature of decoding. This confirms that the bottleneck is largely a function of output token count and model dimension. Therefore, achieving sub-100 ms control loops requires an architectural shift of this Pareto front: future work must focus on distilling these capabilities into significantly smaller backbones (e.g., sub-1B parameters) or employing speculative decoding to decouple latency from message complexity.
@@ V Conclusion and Future Work
    Inference acceleration: speculative multi-token decoding [10] and knowledge distillation [6], hierarchical agent architectures, and lightweight backbones to reconcile fidelity with real-time constraints.
@@
    Extended-context modeling: the context window length of current LLM s is critical for cross-layer reasoning and complex RRC procedures. Longer context techniques [3] or memory-augmented transformers (e.g., RoPE-extended variants [15, 22, 2, 21, 17, 14], ALiBi [14]) will be explored to eliminate segmentation and preserve procedural continuity.